# Gravitational Lensing of Gravitational Waves from Astrophysical Sources: Theory, Detection, and Applications

### 07 May 2026 v1

@@ 4 Detection rates of lensed GW events
where $z_{\ell}$ is the redshift and $dn/d\boldsymbol{x}$ denotes the number density distribution of the intervening objects per unit parameters $\boldsymbol{x}$, such as velocity dispersion $\sigma_{\rm v}$ of elliptical galaxies or halo mass $M_{\rm H}$ of the galaxy clusters. The symbol $S_{\rm cr}$ is the lensing cross-section, defined as the area of the caustics in the source plane, i.e., the area within which the lens Equation (26) can have multiple solutions, equivalently multiple images in the geometric-optics regime. In the wave-optics regime, the cross section can be also defined via the detectability of lensed sources (e.g., Dai et al. 2018; Guo & Lu 2022). With the lensing optical depth, the detection rate $\dot{N}^{\ell}$ can be estimated as
@@ 4.1 Galaxy/cluster-CBC
4.520.34.5-20.3 (14) Dynamical (Zhao & Lu 2021) SIE+external shear ET Chen (2023)
6.629.66.6-29.6 (14) Dynamical (Zhao & Lu 2021) SIE+external shear CE Chen (2023)
@@
Figure 5: Estimates for the Detection rate of strongly lensed CBC and MBBH systems by different GW detectors. The results of CBCs are rescaled by the latest constraint on the local merger rate density from GWTC-4.0, i.e., $R^{\rm E}\sim 19_{-5}^{+7}\rm Gpc^{-3}yr^{-1}$ for sBBHs, $89_{-67}^{+159}\rm Gpc^{-3}yr^{-1}$ for BNSs, and $23_{-13}^{+20}\rm Gpc^{-3}yr^{-1}$ for BHNSs (The LIGO Scientific Collaboration et al. 2025e).
@@
The frequency of GW emitted by CBCs detected by ground-based GW detectors are in the range from ten Hz to several hundred hertz or even higher. Therefore, the galaxy/cluster lensing is in the geometric-optics regime. As for the galaxy-lensed GW signals emitted by CBCs, the detection rate forecasts for various GW detectors and sources have been intensively discussed in the literature (e.g., Ding et al. 2015; Li et al. 2018; Ng et al. 2018; Oguri 2018; Wierda et al. 2021; Yang et al. 2022; Wempe et al. 2022; Chen 2023; Ma et al. 2023b). For a convenient reference for readers, we sort these results in Table 1. Since current predictions on the CBC merger rates still have large uncertainties due to that many complicated dynamical and radiative processes involved in the formation of CBCs are not well understood and the contribution fractions from different mechanisms are difficult to accurately estimate, therefore, it is challenge to estimate the lensed event rate accurately. Recently, the usual approach is to normalize modelled merger‑rate densities to the local rate ($R_{0}$) inferred from LVK observations while preserving the modelled redshift (or mass) evolution. The latest constraints on the local merger rate densities for sBBHs, BNSs, and BHNSs are $R_{0}\sim 19^{+7}_{-5}\rm Gpc^{-3}yr^{-1}$, $\sim 89^{+159}_{-67}\rm Gpc^{-3}yr^{-1}$, and $23^{+20}_{-13}\rm Gpc^{-3}yr^{-1}$, respectively, still with quite large uncertainties (The LIGO Scientific Collaboration et al. 2025e). As seen from Table 1, after rescaling the local merger rate to the same value for each type of CBCs, the predicted values by different authors are different by a factor up to 252-5 for sBBHs, BNSs, and BHNSs, respectively. Such differences are partly due to the different redshift evolution shapes of the adopted merger rate densities, which now cannot be tightly constrained by the LVK observations due to the lack of GW events detected at high redshifts. For example, Li et al. (2018) found that the choice of different SFR density evolution models (consequently different $R(z)$) can introduce a fractional uncertainty of $50\%-200\%$ to $\dot{N}^{\ell}$ depending on adopted GW detectors. In this case, the higher sensitivity of the adopted GW detector, the larger the fractional uncertainty, simply because more distant sources can be detected by the detector where the differences of $R(z)$ between different models are larger. In addition, the choice of the lensing model also affect the estimate of $\dot{N}^{\ell}$. For different lensing models, e.g., SIS, SIE, or EPL model, the estimated lensing cross-section $S_{\rm cr}$ may be different by a factor of $\sim 1-5$, which depending on detail settings of the lensing model, e.g., ellipticity for SIE and power law index for EPL. Furthermore, one may also seen from this table, that the detection rate $\dot{N}^{\ell}$ of galaxy lensed BNS mergers may be substantially lower than that of sBBH merger if assuming 2.5G GW detectors (i.e., aLIGO, LIGO A+ and LIGO Voyager), which is due to their significantly lower S/N at the same redshift. However, when it comes to next generation GW detectors, the detection rate of lensed BNS mergers may be larger than that of sBBH mergers due to their relatively larger merger rate density.
@@
where $N_{\rm pair}=N(N-1)/2$ is the total number of event pairs and $\rm FAP_{pair}$ is the percentage of these pairs with parameter overlaps that mimic the lensed pairs. As shown in Table III of Çalışkan et al. (2023), the combined $\rm FAP_{pair}$ is approximately the order of $10^{-5}\sim 10^{-6}$ for the $95\%$ confidence level lensed GW events detection of the LVK network, which is the combination of false alarm due to mass, sky localization, and coalescence phase overlap. Several works indicated that the third generation GW detectors (ET and CE) will constrain source parameters with precision better by factors of a few tens relative to current LVK measurements, thereby shrinking the overlap region in parameter space by a comparable factor (e.g., Zhao & Wen 2018; Pieroni et al. 2022b). The substantially increased sensitivity of CE and ET is thus expected to reduce the pairwise FAP, $\rm FAP_{pair}$ to the order of $10^{-8}-10^{-10}$ Chen (2023). Then the total FAP for the lensed GW events detected by the third generation detectors is on the order of $10^{-2}-10^{0}$ per year, which is significantly smaller than the predicted number of the detection year rate shown in Table 1. For similar reasons, Barsode et al. (2025) claimed that the improved parameter precision will indeed compensate for the growth in false alarms in the O5 run of LVK, making the first $3\sigma$ detection come true. In addition, the micro-lensing of stars in galaxies or galaxy clusters may also help to more easy identification of lensed GW signals, since it may change the waveform of lensed GW signals differently at different frequencies (e.g, Shan et al. 2024, 2023). Therefore, we are optimistic on the detection of lensed GW events in the upcoming era of the LIGO+, LIGO Voyager and third generation ground-based detectors.
@@ 4.2 Dark matter minihalo/PBH-sBBH
We summarize the estimates for the detection rates of sBBH mergers lensed by dark matter minihalo and PBH in this section. These lens systems are in the wave-optics regime. Figure 7 shows the differential detection rate of sBBH mergers lensed by different dark matter models. As for dark matter minihalo lenses, if adopting the CDM model (NFW density profile for lens) and an S/N difference threshold of $\delta\hat{\rho}_{\rm th}=1$, then the detection rate is about $\dot{N}^{\ell}\sim 6.44\times 10^{-7}$, $4.21\times 10^{-3}$, $0.288$, $1.02$, $9.76$, and $189~\rm yr^{-1}$ for the high-frequency ground-based GW detectors LIGO A+, ET, CE, Gravitational-wave Lunar Observatory for Cosmology (GLOC), and future middle-frequency space-based DECIGO and BBO, respectively (see Guo & Lu 2022). As for the SIDM model, the detection rate is strongly dependent on the size of the core of the lenses. For example, if the core size is rather small, e.g., $r_{1}=0.1$, the detection rate $\dot{N}^{\ell}$ is more or less similar with that of CDM model. However, if we impose a large core, for instance $r_{1}=1.0$, the value of $\dot{N}^{\ell}$ will be substantially suppressed, i.e., to $\sim 2.45\times 10^{-9}$, $8.93\times 10^{-6}$, $4.69\times 10^{-4}$, $2.37$, and $88.2~\rm yr^{-1}$, correspondingly. As for the WDM model, the detection rate is dependent on the particle mass. For example, assuming the particle mass of $30\rm keV$ and $3\rm keV$, the detection rates $\dot{N}^{\ell}$ are about $\sim 0.09$ and $0.002$ per year for CE. Apparently, the middle-frequency GW detectors, such as DECIGO and BBO, have sufficiently high sensitivity to detect lensed GW events via diffractive lensing by minihalos and help to reveal the nature of dark matter, while it is challenging for the current high-frequency ground-based GW detectors to detect such lensing effect. As for fuzzy dark matter model with axion mass of $m_{\rm a}\sim 10^{-17}~\rm eV$, Fairbairn et al. (2023) estimated that the detection rate of lensed events by ET is about $5\times 10^{-4}-4\times 10^{-3}$ and 0.010.070.01-0.07 per year. If the axion mass is substantially low, i.e., $m_{\rm a}\lesssim 10^{-18}~\rm eV$, then the detection of such lensing effect will be extremely difficult. In conclusion, the dark matter nature, either cold, warm, or self-interacting, may be constrained via the detection of the diffractively lensed GW events by future GW detectors.
@@ 4.3 SMBBH-Galaxy
For MBBH merger systems emitting millihertz GWs, the lensing by galaxies is in the geometric-optics regime (e.g., Sereno et al. 2010). By taking galaxies as intervening lens, Gutiérrez & Lagos (2025) and Luo et al. (2025b) predicted that LISA and Taiji can detect about $\sim 0.04-231$ strongly lensed MBBH mergers in a four-year observation period (see also Figure 5), depending on different source populations, including light- and heavy-seed scenarios.
@@ 5 Cosmological Application
The gravitational lensing of GWs holds a wealth of pivotal applications in astrophysics, including the following: (1) Probing the nature of dark matter through distinct GW waveform features and detection rates produced by lensing from intervening dark matter halos with varying intrinsic properties (e.g., Liao et al. 2018, 2020; Cao et al. 2021; Guo & Lu 2022; Cao et al. 2022; Tambalo et al. 2023; Jana et al. 2024; Cheung et al. 2024); (2) Constraining the mass and cosmic abundance of PBHs via lensing signatures in the waveforms of point-like GW sources and the distortion of the stochastic gravitational wave background (Jung & Shin 2019; Diego 2020; Wang et al. 2021b; Urrutia et al. 2023; Sun et al. 2025); (3) Constraining the astrophysical origin and formation channels of sBBH mergers by cross-correlating lensed GW events with their host galaxies (e.g., Chen et al. 2022b; He et al. 2022; Chen 2023); (4) Placing tight constraint on the GW propagation speed by comparing the arrival times and time-delays between lensed GWs and their EM counterparts (Fan et al. 2017; Collett & Bacon 2017); (5) Setting limits on the graviton mass and probing the nature of gravity via the time-delays and waveform-morphology modifications induced by gravitational lensing (e.g., Yang et al. 2019; Liao et al. 2019; Mukherjee et al. 2020; Chung & Li 2021).
Furthermore, the gravitational lensing of GWs can also be applied as an independent probe to constrain cosmological parameters (e.g., Liao et al. 2017; Li et al. 2019b; Hou et al. 2020, 2021a; Hannuksela et al. 2020; Chen et al. 2025a), especially it may be helpful in resolving the Hubble tension. In this section, we outline two methods to infer the cosmological parameters, namely “Bright Lensed Siren” (Liao et al. 2017) and “Dark Lensed Siren” (Chen et al. 2025a). In general, both methods take the advantage of the extremely precise measurement on the time-delay $\Delta t_{i,j}$ among lensed images of GW signals,
@@ 5.1 Bright Lensed Siren
The “Bright Lensed Siren” was first proposed by analogizing to the time-delay cosmography method of strongly-lensed Quasars (e.g., Wong et al. 2020). In this method, one may reconstruct the map of $\Delta\phi$ from observations of lensed host galaxies and then determine the specific $\Delta\phi_{i,j}$ for the $i$-th and $j$-th images of the lensed GW signals by the positions of lensed EM counterparts, i.e., kilonovae and afterglows, in the host galaxies. The advantage of this method compared with the traditional time-delay cosmography method lies within the following two aspects. First, the time resolution of GW detectors can be extremely precise, normally at the scale of several ms, which is negligible comparing with time-delays between different images (typically days to months). However, the relative uncertainty of time-delay measurement for lensed Quasars is normally on the order of $\sim 3-5\%$. Second, comparing with the case of lensed Quasars, the potential map $\Delta\phi$ can be reconstructed more precisely as there is no contamination from AGN light. By performing Monte Carlo Simulations, Liao et al. (2017) found that using about $\sim 10$ such lensed GW+EM systems can provide a $\sim 0.68\%$ uncertainty of the Hubble constant $H_{0}$ with next-generation GW detectors.
Note that in the above “Bright Lensed Siren” method requires observed EM counterparts of the lensed events to locate the GW source in its host galaxies (e.g., Liu & Liao 2025) and thereby measure the specific potential difference $\Delta\phi_{i,j}$ between images. Consequently, this approach applies primarily to BNS and BHNS systems, since sBBH mergers typically lack observable EM counterparts, unless they occur in special environments (e.g., AGN disks). The method’s effectiveness therefore depends strongly on the joint GW-EM detection rate. Ma et al. (2023b) and the follow-up studies by Chen et al. (2025) outlined searching strategies for lensed multi-messenger BNS signals and emphasized the importance of identifying lensed host galaxies. By this strategy, the joint GW-EM detection rate for lensed systems can be estimated as
@@
Here, $\dot{N}^{\ell}_{\rm GW}$ denotes the rate of lensed GW events, which is estimated to be on the order of one to several tens per year for the third-generation GW detectors such as CE and ET (see Table 1); $f_{\rm EM}$ represents the fraction of lensed GW events that have detectable lensed EM counterparts, which is set largely by the limiting flux (or magnitude) of a telescope; for example, adopting the infrared band F158 of the Roman Space Telescope (RST), yields $f_{\rm EM}\sim 0.1$ for kilonovae and $\sim 0.01$ for afterglows; $f_{\rm Host}$ denotes the fraction of host galaxies that can be identified as lensed systems in a given survey over a certain sky area, which typically ranges from $\sim 20\%$ to $50\%$, depending on the telescope’s limiting magnitude and the criteria adopted for lens identification; $P_{\rm Host}$ represents the probability of successfully matching the associated lensed host galaxies, depending directly on the localization uncertainty, and $P_{\rm Host}\sim 0.1$ for a localization precision of $1$ deg2 and $\sim 1$ for $0.1$ deg2 if assuming a random matching strategy. By detailed modeling of the EM population of BNS mergers, utilizing both the results from observation and numerical simulations to estimate those terms above, Chen et al. (2025) found that the joint detection of lensed GW with kilonova is promising for the Roman Space Telescope (RST) and the detection rate $\dot{N}^{\ell}_{\rm GW+K}$ is about $\sim 0.45_{-0.34}^{+0.80}$, $0.55_{-0.41}^{+0.98}$, and $0.06_{-0.05}^{+0.11}$ yr-1 with the RST F106, F158, F213 bands, respectively. As for other telescopes such as CSST and Euclid, the expectation detection rate may be $\sim 10$ times smaller, limited by their small limiting magnitude (Ma et al. 2023b). The joint detection of lensed afterglow signals in the optical and radio bands is however even more difficult, even with the RST and Square Kilometer Array (SKA). Therefore, achieving a $1\%$ precision measurement of the Hubble constant may require waiting 102010-20 yrs after third generation GW detectors come online.

### 16 May 2026 v2

@@ 4 Detection rates of lensed GW events
where $z_{\ell}$ is the redshift and $dn/d\boldsymbol{x}$ denotes the number density distribution of the intervening objects per unit parameters $\boldsymbol{x}$, such as velocity dispersion $\sigma_{\rm v}$ of elliptical galaxies or halo mass $M_{\rm H}$ of the galaxy clusters. The symbol $S_{\rm cr}$ is the lensing cross-section, defined as the area of the caustics in the source plane, i.e., the area within which the lens Equation (26) can have multiple solutions, equivalently multiple images in the geometric-optics regime. In the wave-optics regime, the cross section can be also defined via the detectability of lensed sources (e.g., Dai et al. 2018; Guo and Lu 2022). With the lensing optical depth, the detection rate $\dot{N}^{\ell}$ can be estimated as
@@ 4.1 Galaxy/cluster-CBC
4.520.34.5-20.3 (14) Dynamical (Zhao and Lu 2021) SIE+external shear ET Chen (2023)
6.629.66.6-29.6 (14) Dynamical (Zhao and Lu 2021) SIE+external shear CE Chen (2023)
@@
Figure 5: Estimates for the Detection rate of strongly lensed CBC and MBBH systems by different GW detectors. The results of CBCs are rescaled by the latest constraint on the local merger rate density from GWTC-4.0, i.e., $R^{\rm E}\sim 19_{-5}^{+7}\rm Gpc^{-3}yr^{-1}$ for sBBHs, $89_{-67}^{+159}\rm Gpc^{-3}yr^{-1}$ for BNSs, and $23_{-13}^{+20}\rm Gpc^{-3}yr^{-1}$ for BHNSs (The LIGO Scientific Collaboration et al. 2025b).
@@
The frequency of GW emitted by CBCs detected by ground-based GW detectors are in the range from ten Hz to several hundred hertz or even higher. Therefore, the galaxy/cluster lensing is in the geometric-optics regime. As for the galaxy-lensed GW signals emitted by CBCs, the detection rate forecasts for various GW detectors and sources have been intensively discussed in the literature (e.g., Ding et al. 2015; Li et al. 2018; Ng et al. 2018; Oguri 2018; Wierda et al. 2021; Yang et al. 2022; Wempe et al. 2022; Chen 2023; Ma et al. 2023b). For a convenient reference for readers, we sort these results in Table 1. Since current predictions on the CBC merger rates still have large uncertainties due to that many complicated dynamical and radiative processes involved in the formation of CBCs are not well understood and the contribution fractions from different mechanisms are difficult to accurately estimate, therefore, it is challenge to estimate the lensed event rate accurately. Recently, the usual approach is to normalize modelled merger‑rate densities to the local rate ($R_{0}$) inferred from LVK observations while preserving the modelled redshift (or mass) evolution. The latest constraints on the local merger rate densities for sBBHs, BNSs, and BHNSs are $R_{0}\sim 19^{+7}_{-5}\rm Gpc^{-3}yr^{-1}$, $\sim 89^{+159}_{-67}\rm Gpc^{-3}yr^{-1}$, and $23^{+20}_{-13}\rm Gpc^{-3}yr^{-1}$, respectively, still with quite large uncertainties (The LIGO Scientific Collaboration et al. 2025b). As seen from Table 1, after rescaling the local merger rate to the same value for each type of CBCs, the predicted values by different authors are different by a factor up to 252-5 for sBBHs, BNSs, and BHNSs, respectively. Such differences are partly due to the different redshift evolution shapes of the adopted merger rate densities, which now cannot be tightly constrained by the LVK observations due to the lack of GW events detected at high redshifts. For example, Li et al. (2018) found that the choice of different SFR density evolution models (consequently different $R(z)$) can introduce a fractional uncertainty of $50\%-200\%$ to $\dot{N}^{\ell}$ depending on adopted GW detectors. In this case, the higher sensitivity of the adopted GW detector, the larger the fractional uncertainty, simply because more distant sources can be detected by the detector where the differences of $R(z)$ between different models are larger. In addition, the choice of the lensing model also affect the estimate of $\dot{N}^{\ell}$. For different lensing models, e.g., SIS, SIE, or EPL model, the estimated lensing cross-section $S_{\rm cr}$ may be different by a factor of $\sim 1-5$, which depending on detail settings of the lensing model, e.g., ellipticity for SIE and power law index for EPL. Furthermore, one may also seen from this table, that the detection rate $\dot{N}^{\ell}$ of galaxy lensed BNS mergers may be substantially lower than that of sBBH merger if assuming 2.5G GW detectors (i.e., aLIGO, LIGO A+ and LIGO Voyager), which is due to their significantly lower S/N at the same redshift. However, when it comes to next generation GW detectors, the detection rate of lensed BNS mergers may be larger than that of sBBH mergers due to their relatively larger merger rate density.
@@
where $N_{\rm pair}=N(N-1)/2$ is the total number of event pairs and $\rm FAP_{pair}$ is the percentage of these pairs with parameter overlaps that mimic the lensed pairs. As shown in Table III of Çalışkan et al. (2023), the combined $\rm FAP_{pair}$ is approximately the order of $10^{-5}\sim 10^{-6}$ for the $95\%$ confidence level lensed GW events detection of the LVK network, which is the combination of false alarm due to mass, sky localization, and coalescence phase overlap. Several works indicated that the third generation GW detectors (ET and CE) will constrain source parameters with precision better by factors of a few tens relative to current LVK measurements, thereby shrinking the overlap region in parameter space by a comparable factor (e.g., Zhao and Wen 2018; Pieroni et al. 2022b). The substantially increased sensitivity of CE and ET is thus expected to reduce the pairwise FAP, $\rm FAP_{pair}$ to the order of $10^{-8}-10^{-10}$ Chen (2023). Then the total FAP for the lensed GW events detected by the third generation detectors is on the order of $10^{-2}-10^{0}$ per year, which is significantly smaller than the predicted number of the detection year rate shown in Table 1. For similar reasons, Barsode et al. (2025a) claimed that the improved parameter precision will indeed compensate for the growth in false alarms in the O5 run of LVK, making the first $3\sigma$ detection come true. In addition, the micro-lensing of stars in galaxies or galaxy clusters may also help to more easy identification of lensed GW signals, since it may change the waveform of lensed GW signals differently at different frequencies (e.g, Shan et al. 2024, 2023). Therefore, we are optimistic on the detection of lensed GW events in the upcoming era of the LIGO+, LIGO Voyager and third generation ground-based detectors.
@@ 4.2 Dark matter minihalo/PBH-sBBH
We summarize the estimates for the detection rates of sBBH mergers lensed by dark matter minihalo and PBH in this section. These lens systems are in the wave-optics regime. Figure 7 shows the differential detection rate of sBBH mergers lensed by different dark matter models. As for dark matter minihalo lenses, if adopting the CDM model (NFW density profile for lens) and an S/N difference threshold of $\delta\hat{\rho}_{\rm th}=1$, then the detection rate is about $\dot{N}^{\ell}\sim 6.44\times 10^{-7}$, $4.21\times 10^{-3}$, $0.288$, $1.02$, $9.76$, and $189~\rm yr^{-1}$ for the high-frequency ground-based GW detectors LIGO A+, ET, CE, Gravitational-wave Lunar Observatory for Cosmology (GLOC), and future middle-frequency space-based DECIGO and BBO, respectively (see Guo and Lu 2022). As for the SIDM model, the detection rate is strongly dependent on the size of the core of the lenses. For example, if the core size is rather small, e.g., $r_{1}=0.1$, the detection rate $\dot{N}^{\ell}$ is more or less similar with that of CDM model. However, if we impose a large core, for instance $r_{1}=1.0$, the value of $\dot{N}^{\ell}$ will be substantially suppressed, i.e., to $\sim 2.45\times 10^{-9}$, $8.93\times 10^{-6}$, $4.69\times 10^{-4}$, $2.37$, and $88.2~\rm yr^{-1}$, correspondingly. As for the WDM model, the detection rate is dependent on the particle mass. For example, assuming the particle mass of $30\rm keV$ and $3\rm keV$, the detection rates $\dot{N}^{\ell}$ are about $\sim 0.09$ and $0.002$ per year for CE. Apparently, the middle-frequency GW detectors, such as DECIGO and BBO, have sufficiently high sensitivity to detect lensed GW events via diffractive lensing by minihalos and help to reveal the nature of dark matter, while it is challenging for the current high-frequency ground-based GW detectors to detect such lensing effect. As for fuzzy dark matter model with axion mass of $m_{\rm a}\sim 10^{-17}~\rm eV$, Fairbairn et al. (2023) estimated that the detection rate of lensed events by ET is about $5\times 10^{-4}-4\times 10^{-3}$ and 0.010.070.01-0.07 per year. If the axion mass is substantially low, i.e., $m_{\rm a}\lesssim 10^{-18}~\rm eV$, then the detection of such lensing effect will be extremely difficult. In conclusion, the dark matter nature, either cold, warm, or self-interacting, may be constrained via the detection of the diffractively lensed GW events by future GW detectors.
@@ 4.3 MBBH-Galaxy
On the one hand, for MBBH merger systems emitting millihertz GWs, the lensing by galaxies is in the geometric-optics regime (e.g., Sereno et al. 2010). By taking galaxies as intervening lens, Gutiérrez and Lagos (2025) and Luo,Yuchao et al. (2025) predicted that LISA and Taiji can detect about $\sim 0.04-231$ strongly lensed MBBH mergers in a four-year observation period (see also Figure 5), depending on different source populations, including light- and heavy-seed scenarios. On the other hand, MBBH mergers may also be lensed by low-mass DM halos or subhalos, which belongs to the wave-optics regime for their smaller sizes (Çalı s , kan et al. 2023b). By considering the diffraction waveform including merger, ring-down and higher terms, Çalı s , kan et al. (2023a) systematically estimate the detection rate of MBBH mergers with diffractive lensing signatures for LISA, which is $\sim 0.01-7.96$ for $1\sigma$ detection threshold. Notably, these rates are strongly subject to the DM HMF, such as the theory-based Press-Schechter halo mass function and an observation-based model derived from Sloan Digital Sky Survey they adopted and similarly also MBBH populations.
@@ 5 Cosmological Application
The gravitational lensing of GWs holds a wealth of pivotal applications in astrophysics, including the following: (1) Probing the nature of dark matter through distinct GW waveform features and detection rates produced by lensing from intervening dark matter halos with varying intrinsic properties (e.g., Liao et al. 2018, 2020; Cao et al. 2021; Guo and Lu 2022; Basak et al. 2022; Cao et al. 2022; Tambalo et al. 2023; Jana et al. 2024; Cheung et al. 2024; Barsode et al. 2024); (2) Constraining the mass and cosmic abundance of PBHs via lensing signatures in the waveforms of point-like GW sources and the distortion of the stochastic gravitational wave background (Jung and Shin 2019; Diego 2020; Wang et al. 2021b; Urrutia et al. 2023; Sun et al. 2025); (3) Constraining the astrophysical origin and formation channels of sBBH mergers by cross-correlating lensed GW events with their host galaxies (e.g., Chen et al. 2022b; He et al. 2022; Chen 2023); (4) Placing tight constraint on the GW propagation speed by comparing the arrival times and time-delays between lensed GWs and their EM counterparts (Fan et al. 2017; Collett and Bacon 2017); (5) Setting limits on the graviton mass and probing the nature of gravity via the time-delays and waveform-morphology modifications induced by gravitational lensing (e.g., Yang et al. 2019; Liao et al. 2019; Mukherjee et al. 2020; Chung and Li 2021; Wright et al. 2025).
Furthermore, the gravitational lensing of GWs can also be applied as an independent probe to constrain cosmological parameters (e.g., Liao et al. 2017b; Li et al. 2019b; Hou et al. 2020, 2021a; Hannuksela et al. 2020; Jana et al. 2023; Maity et al. 2025; Chen et al. 2025b), especially it may be helpful in resolving the Hubble tension. In this section, we outline two methods to infer the cosmological parameters, namely “Bright Lensed Siren” (Liao et al. 2017a) and “Dark Lensed Siren” (Chen et al. 2025b). In general, both methods take the advantage of the extremely precise measurement on the time-delay $\Delta t_{i,j}$ among lensed images of GW signals,
@@ 5.1 Bright Lensed Siren
The “Bright Lensed Siren” was first proposed by analogizing to the time-delay cosmography method of strongly-lensed Quasars (e.g., Wong et al. 2020). In this method, one may reconstruct the map of $\Delta\phi$ from observations of lensed host galaxies and then determine the specific $\Delta\phi_{i,j}$ for the $i$-th and $j$-th images of the lensed GW signals by the positions of lensed EM counterparts, i.e., kilonovae and afterglows, in the host galaxies. The advantage of this method compared with the traditional time-delay cosmography method lies within the following two aspects. First, the time resolution of GW detectors can be extremely precise, normally at the scale of several ms, which is negligible comparing with time-delays between different images (typically days to months). However, the relative uncertainty of time-delay measurement for lensed Quasars is normally on the order of $\sim 3-5\%$. Second, comparing with the case of lensed Quasars, the potential map $\Delta\phi$ can be reconstructed more precisely as there is no contamination from AGN light. By performing Monte Carlo Simulations, Liao et al. (2017a) found that using about $\sim 10$ such lensed GW+EM systems can provide a $\sim 0.68\%$ uncertainty of the Hubble constant $H_{0}$ with next-generation GW detectors.
Note that in the above “Bright Lensed Siren” method requires observed EM counterparts of the lensed events to locate the GW source in its host galaxies (e.g., Liu and Liao 2025) and thereby measure the specific potential difference $\Delta\phi_{i,j}$ between images. Consequently, this approach applies primarily to BNS and BHNS systems, since sBBH mergers typically lack observable EM counterparts, unless they occur in special environments (e.g., AGN disks). The method’s effectiveness therefore depends strongly on the joint GW-EM detection rate. Ma et al. (2023b) and the follow-up studies by Chen et al. (2025a) outlined searching strategies for lensed multi-messenger BNS signals and emphasized the importance of identifying lensed host galaxies. By this strategy, the joint GW-EM detection rate for lensed systems can be estimated as
@@
Here, $\dot{N}^{\ell}_{\rm GW}$ denotes the rate of lensed GW events, which is estimated to be on the order of one to several tens per year for the third-generation GW detectors such as CE and ET (see Table 1); $f_{\rm EM}$ represents the fraction of lensed GW events that have detectable lensed EM counterparts, which is set largely by the limiting flux (or magnitude) of a telescope; for example, adopting the infrared band F158 of the Roman Space Telescope (RST), yields $f_{\rm EM}\sim 0.1$ for kilonovae and $\sim 0.01$ for afterglows; $f_{\rm Host}$ denotes the fraction of host galaxies that can be identified as lensed systems in a given survey over a certain sky area, which typically ranges from $\sim 20\%$ to $50\%$, depending on the telescope’s limiting magnitude and the criteria adopted for lens identification; $P_{\rm Host}$ represents the probability of successfully matching the associated lensed host galaxies, depending directly on the localization uncertainty, and $P_{\rm Host}\sim 0.1$ for a localization precision of $1$ deg2 and $\sim 1$ for $0.1$ deg2 if assuming a random matching strategy. By detailed modeling of the EM population of BNS mergers, utilizing both the results from observation and numerical simulations to estimate those terms above, Chen et al. (2025a) found that the joint detection of lensed GW with kilonova is promising for the Roman Space Telescope (RST) and the detection rate $\dot{N}^{\ell}_{\rm GW+K}$ is about $\sim 0.45_{-0.34}^{+0.80}$, $0.55_{-0.41}^{+0.98}$, and $0.06_{-0.05}^{+0.11}$ yr-1 with the RST F106, F158, F213 bands, respectively. As for other telescopes such as CSST and Euclid, the expectation detection rate may be $\sim 10$ times smaller, limited by their small limiting magnitude (Ma et al. 2023b). The joint detection of lensed afterglow signals in the optical and radio bands is however even more difficult, even with the RST and Square Kilometer Array (SKA). Therefore, achieving a $1\%$ precision measurement of the Hubble constant may require waiting 102010-20 yrs after third generation GW detectors come online.